OpenStudy (anonymous):

Help! Urgent! Will medal! Solve the inequality:

OpenStudy (anonymous):

\[4x+15\le 19\]

OpenStudy (lxelle):

4x+15≤19 you minus 15 on both sides leaving only 4x≤19-15, which is 4x≤4 solve from there.

OpenStudy (anonymous):

What? That kinda confused me

OpenStudy (igreen):

\(4x + 15 \le 19\) Subtract 15 to both sides. \(4x \le 4\) Now divide. What's 4 / 4? @pigglet

OpenStudy (anonymous):

1

OpenStudy (igreen):

Yep, so that gives us: \(x \le 1\)

OpenStudy (anonymous):

thank you!

OpenStudy (igreen):

No problem.